OpenStudy (anonymous):

graph y=-3/4x (DON'T GRAPH I HAVE A QUESTION)

OpenStudy (anonymous):

If \[y=-\frac{ 3 }{ 4 }\] because of the negative does it make 4 negative as well or just the three? For example if I were graphing this and used rise over run

OpenStudy (anonymous):

I think I understand what you're trying to ask... y = -3/4x. You're asking what that negative refers to. Remember that a negative simply means you're multiplying by (-1). So y = (-1/1) * (3/4). Y can ALSO equal (1/-1) * (3/4), because we know that 1/-1 equals -1, and -1/1 equals -1. So our two equations from y= (-)(3/4) are: y = -3/4 and y= 3/-4. Graph by using the rise over run method. You'll see that they are the same!

OpenStudy (anonymous):

I see, thank you for explaining:)